# REG4Rec: Reasoning-Enhanced Generative Model for Large-Scale Recommendation Systems

### v3

@@ Abstract.
Sequential recommendation aims to predict a user’s next action in large-scale recommender systems. While traditional methods often suffer from insufficient information interaction, recent generative recommendation models partially address this issue by directly generating item predictions. To better capture user intents, recent studies have introduced a reasoning process into generative recommendation, significantly improving recommendation performance. However, these approaches are constrained by the singularity of item semantic representations, facing challenges such as limited diversity in reasoning pathways and insufficient reliability in the reasoning process. To tackle these issues, we introduce REG4Rec, a reasoning-enhanced generative model that constructs multiple dynamic semantic reasoning paths alongside a self-reflection process, ensuring high-confidence recommendations. Specifically, REG4Rec utilizes an MoE-based parallel quantization codebook (MPQ) to generate multiple unordered semantic tokens for each item, thereby constructing a larger-scale diverse reasoning space. Furthermore, to enhance the reliability of reasoning, we propose a training reasoning enhancement stage, which includes Preference Alignment for Reasoning (PARS) and a Multi-Step Reward Augmentation (MSRA) strategy. PARS uses reward functions tailored for recommendation to enhance reasoning and reflection, while MSRA introduces future multi-step actions to improve overall generalization. During inference, we further incorporate Consistency-Oriented Self-Reflection for Pruning (CORP) to correct or discard inconsistent reasoning paths, thereby mitigating noise and preventing the propagation of erroneous reasoning. Lastly, we develop an efficient offline training strategy for large-scale recommendation. Experiments on actual datasets have verified that REG4Rec has achieved outstanding performance, realizing a performance gain of up to 16.59%. Online experiments demonstrate its substantial practical value.
Generative Recommendation, Reasoning Process, Self-Reflection
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Retrieval models and ranking
1. Introduction
@@
Sequential recommendation (SR), due to its ability to predict users’ next actions and uncover latent intentions by analyzing historical behavior patterns (Boka et al., 2024), is widely applied to platforms such as e-commerce websites (Deng et al., 2025b; Lin et al., 2024), video streaming services (Wang et al., 2025), and news portals (Li et al., 2022). Recently, generative models have achieved significant breakthroughs in sequential recommendation (Zhai et al., 2024). Generative Recommendation (GR) replaces traditional numeric item IDs with fixed, unique sequences of semantic tokens for item representation. By leveraging the semantic token’s content representation capability and generative models’ intent understanding capacity, GR dynamically generates items consistent with users’ latent preferences and has quickly emerged as a new paradigm (Rajput et al., 2023; Yang et al., 2025).
@@
Recently, several studies have integrated intermediate reasoning process into GR (Figure 1(a)), generating item semantic tokens to construct paths toward final predictions and achieving significant performance gains (Yang et al., 2025; Tang et al., 2025; Zhang et al., 2025b). However, existing models face several limitations. First, each item is deterministically encoded into a fixed sequence of semantic tokens, producing rigid representations that conflict with the inherent uncertainty of reasoning. Since inference paths are constructed from these fixed item representations, the singularity of the representations directly results in a lack of diversity in the exploration space of reasoning pathways. Items have multi-dimensional attributes such as brand, style, and color, while different users may have diverse reasons for interest in the same item. Consequently, the constrained reasoning space of existing models fails to adequately capture this diversity of user intents, ultimately leading to suboptimal recommendation performance. Second, current methods lack mechanisms to assess or calibrate the reliability of reasoning paths, which may introduce uncontrolled noise during the reasoning process and reduce recommendation accuracy. To overcome these limitations, we aim to develop a reliable GR reasoning framework that supports diverse reasoning paths. To this end, we propose two research problems: (RQ1) How can we design flexible reasoning paths to accommodate diverse user intents? (RQ2) How can we ensure a reliable reasoning process to enhance recommendation quality?
@@
Figure 1. Comparison of reasoning in GR: (a) GR with a limited reasoning process; (b) our REG4Rec enables diverse and reliable reasoning process to better capture the variety of users’ intents.
$x$.
@@
To address these questions, we draw on core mechanisms of Large Language Models (LLMs). Their success in complex tasks stems from two advances: adopting dynamic, multi-step reasoning beyond linear thought, and incorporating self-reflective critique to improve output reliability (Yao et al., 2023b, a; Guo et al., 2025). Building on these ideas, we introduce REG4Rec, a reasoning-enhanced generative recommendation model (Figure 1 (b)). To design flexible reasoning paths (RQ1), we introduce the Mixture-of-Experts (MoE)–based Parallel Quantization Codebooks (MPQ) for each item. Given multimodal inputs, MPQ uses parallel autoencoder experts to generate distinct, unordered semantic codebooks, each representing a different facet of the item. At each step, REG4Rec generates parallel predictions from all codebooks and dynamically selects the token with the highest confidence, enabling a vast combinatorial variety of paths. To ensure a reliable reasoning process (RQ2), we propose a comprehensive framework for training and inference. Specifically, we design a Preference Alignment Reward for Selection (PARS) using reinforcement learning (RL) (Shao et al., 2024) to guide the model in selecting inference paths with high consistency and confidence. Furthermore, we propose Multi-Step Reward Augmentation (MSRA), which extends the reward horizon to future user actions, enabling better capture of long-term preferences and reducing noise from stochastic user behavior. During inference, we employ Consistency-Oriented Self-Reflection for Pruning (CORP) to identify and correct or discard inconsistent reasoning paths, thereby mitigating noise and preventing the propagation of erroneous reasoning.
@@
We conduct extensive experiments on four datasets (three public datasets and a large-scale industrial dataset from an online advertising system), comparing REG4Rec with six state-of-the-art baselines, such as GR models with reasoning process like ReaRec and STREAM. REG4Rec consistently outperforms all baselines across every scenario, achieving performance gains of up to 16.59%. To validate its real-world effectiveness, we deploy REG4Rec on a large-scale commercial advertising platform. The subsequent online A/B test yielded a 5.60% increase in Advertising Revenue, a 1.81% increase in Click-Through Rate (CTR), and a 3.29% increase in Gross Merchandise Volume (GMV). These results demonstrate the significant practical value and real-world impact of our model.
@@
2. Related Work
2.1. Sequential Recommendation
SR is a fundamental task that predicts a user’s next action from their interaction history (Yang et al., 2025; Deng et al., 2025a, b). While early methods employed CNNs (Chen et al., 2022; Yan et al., 2019) and RNNs (Hidasi et al., 2015; Hidasi and Karatzoglou, 2018), recent work has centered on transformer-based architectures. For instance, SASRec (Kang and McAuley, 2018) frames SR as an autoregressive prediction problem, BERT4Rec (Sun et al., 2019) employs bidirectional self-attention to leverage both preceding and succeeding contexts. With billions of items, this approach exacerbates the Matthew effect (Liu and Huang, 2021) and suffers from data sparsity, leading to poorly trained embeddings and limited scalability (Zhai et al., 2024). To mitigate this, GR models have emerged, replacing vast item ID vocabularies with compact, learned semantic tokens. A prominent technique for this is the Residual Quantized VAE (RQ-VAE), used by models like TIGER (Rajput et al., 2023), which encodes items into a hierarchical sequence of discrete tokens. This approach, also seen in models like LIGER (Yang et al., 2024) and COBRA (Yang et al., 2025), allows for generation within a compressed latent space.
@@ 2.1. Sequential Recommendation
Existing GR models, which generate recommendations in a single step, often overlook the reasoning needed to capture complex user intents. In contrast, we propose a reasoning-enhanced model that adaptively selects reliable reasoning paths, delivering more accurate and insightful recommendations.
$2.$ Reasoning-Enhanced Generation
@@
Reasoning is fundamental to solving complex problems (Manning, 2022; Wu et al., 2024). Recent breakthroughs in LLMs have moved beyond single-step generation, driven by two key advancements. First, LLMs create adaptive, structured reasoning flows. Instead of following a single linear path, models now explore a dynamic space of possibilities, either internally through branching thought-paths (e.g., Tree of Thoughts (Yao et al., 2023a)) or externally by interacting with tools (e.g., ReAct (Yao et al., 2023b)). Second, they integrate reflection mechanisms for self-critique (Guo et al., 2025; Jaech et al., 2024). This involves a model drafting, evaluating, and iteratively correcting its own solutions, which significantly enhances output reliability. Inspired by these LLM advancements, researchers have begun to incorporate reasoning into recommendation systems (Tang et al., 2025; Zhang et al., 2025b). However, current approaches often implement these principles in a limited fashion. For instance, while ReaRec (Tang et al., 2025) generates intermediate outcomes, its implicit nature lacks the verifiable structure required for robust reasoning. Similarly, although STREAM (Zhang et al., 2025b) constructs explicit reasoning paths, its approach is overly rigid: it confines paths to textually similar items and represents each item with a fixed sequence of semantic tokens. This rigidity prevents the model from capturing the diverse intents users may have for the same item, which necessitates more adaptive reasoning.
@@
Figure 2. REG4Rec introduces MPQ for generating flexible and reliable reasoning paths. During training, CRSS diversifies path combinations, while PARS selects consistent, confident paths. MSRA extends the reward horizon to capture long-term preferences and reduce noise caused by stochastic behavior. During inference, CORP prunes inconsistent paths, preventing the propagation of errors in the reasoning process.
$x$.
3. Methodology
3.1. Problem Formulation
This section formalizes the GR task. Let $\mathcal{I}$ denote the set of all items. For a given user $u$, the historical interaction sequence is defined as $\mathcal{S}_{u}=(i_{1},i_{2},\ldots,i_{T})$, where $i_{t}\in\mathcal{I}$ is the item interacted with at time step $t$. The primary objective is to predict the user’s next interaction, $i_{T+1}\in\mathcal{I}$. Inspired by prior work (Rajput et al., 2023; Yang et al., 2025), we integrate the GR model into the retrieval stage of recommender systems. A common approach is to build an encoder-decoder GR model. The model takes $\mathcal{S}_{u}$ as input and estimates a distribution $p(i|\mathcal{S}_{u})$ over $i\in\mathcal{I}$. During inference, the model retrieves the top-NN items by sorting $p(i|\mathcal{S}_{u})$ in descending order.
3.2. Overview of REG4Rec
@@
Our method addresses two key objectives: (i) designing flexible reasoning steps to accommodate diverse user intents (Section 3.3); and (ii) enhancing reasoning reliability to improve recommendation performance (Sections 3.4 and 3.5).
$3.$ Flexible Reasoning Paths
@@
3.3.1. MPQ: Moe-based Parallel Quantization Codebooks
At the item tokenization stage, in contrast to traditional RQ codebooks (Yang et al., 2025), we propose the MPQ for the reasoning process in REG4Rec. MPQ uses multiple codebooks in parallel, which are order-invariant. Each codebook captures a distinct semantic dimension of an item, yielding a factorized representation without a predefined hierarchy. Leveraging an MoE-style gating mechanism (Ma et al., 2018), MPQ enables information sharing across codebooks and adaptively adjusts their granularity and diversity to better capture semantics. As a result, MPQ balances the information used at each reasoning step with the overall size of the reasoning space.
Concretely, for an item $i$, its multi-modal features (e.g., textual, visual, and behavioral) are extracted with a pretrained model (Xing et al., 2025) and concatenated into a unified feature vector $v_{i}$. Let $M$ denote the number of parallel autoencoder experts $\{E_{r}\}^{M}_{r=1}$. Each expert $E_{r}$ is associated with a codebook $\mathcal{C}^{r}=\{z^{r}_{1},z^{r}_{2},\ldots,z^{r}_{D}\}$, where $z^{r}_{j}$ is a codeword and $D$ is the codebook size. Each expert operates independently, takes $v_{i}$ as input and projects it into a latent spaces $e^{r}_{i}=E_{r}(v_{i})$. The experts then select the most similar codeword in $\mathcal{C}_{r}$ to $e^{r}_{i}$, using its index to represent a semantic token for item ii. Formally, the index $c^{i}_{r}$ which is selected by the $r$-th expert represent the semantic token, is calculated as:
(1) $c^{i}_{r}=\arg\min_{j}\|e^{r}_{i}-z^{r}_{j}\|_{2},$
where $\arg\min$ returns the index of the codeword with the smallest L2 distance (Boyd and Vandenberghe, 2004) to $e^{r}_{i}$. After obtaining $z^{r}_{i}$ from each expert, we employ a routing gate to aggregate the semantic contributions from all codebooks. The final encoding vector is $\mathbf{q}_{i}=\sum_{r=1}^{M}h^{r}(v_{i})z^{r}_{i}$, where $h^{r}(v_{i})$ denotes the routing weight assigned to expert $r$, based on the input feature vector $v_{i}$.

### v4

@@ Abstract
In large-scale e-commerce platforms, sequential recommendation must address diverse user behaviors driven by distinct shopping intents. Users may engage in routine purchases, flash sale shopping, cross-category exploration, or long-term decision, each requiring tailored interaction understanding. While traditional methods struggle with these complexities, generative models offer improved prediction capabilities. To better capture intent diversity in e-commerce scenarios, recent studies have introduced a reasoning process into generative recommendation. However, these approaches are constrained by the singularity of item semantic representations, facing challenges such as limited diversity in reasoning pathways and insufficient reliability in the reasoning process. To tackle these issues, we introduce REG4Rec, a reasoning-enhanced generative model that constructs multiple dynamic semantic reasoning paths alongside a self-reflection process, ensuring high-confidence recommendations. Specifically, REG4Rec utilizes an MoE-based parallel quantization codebook (MPQ) to generate multiple unordered semantic tokens for each item, thereby constructing a larger-scale diverse reasoning space. Furthermore, to enhance the reliability of reasoning, we propose a training reasoning enhancement stage, which includes Preference Alignment for Reasoning (PARS) and a Multi-Step Reward Augmentation (MSRA) strategy. PARS uses reward functions tailored for recommendation to enhance reasoning and reflection, while MSRA introduces future multi-step actions to improve overall generalization. During inference, we further incorporate Consistency-Oriented Self-Reflection for Pruning (CORP) to correct or discard inconsistent reasoning paths, thereby mitigating noise and preventing the propagation of erroneous reasoning. Lastly, we develop an efficient offline training strategy for large-scale recommendation.
Lastly, we develop an efficient offline training strategy tailored for large‑scale recommendation pipelines. This strategy combines dynamic embedding data management, compiler, and a layer‑adaptive dynamic quantization controller (LADQ) to reduce training time and overall hardware cost while preserving model accuracy for post‑training refinement. Experiments on actual datasets have verified that REG4Rec has achieved outstanding performance, realizing a performance gain of up to 16.59%. Online experiments demonstrate its substantial practical value, including a live deployment and A/B test on Alibaba Group’s online advertising platform.
I Introduction
@@
Sequential recommendation (SR), due to its ability to predict users’ next actions and uncover latent intentions by analyzing historical behavior patterns [12], is widely applied to platforms such as e-commerce websites [5, 11], video streaming services [9], and news portals [8]. Recently, generative models have achieved significant breakthroughs in sequential recommendation [13]. Generative Recommendation (GR) replaces traditional numeric item IDs with fixed, unique sequences of semantic tokens for item representation. By leveraging the semantic token’s content representation capability and generative models’ intent understanding capacity, GR dynamically generates items consistent with users’ latent preferences and has quickly emerged as a new paradigm [18, 10].
@@
Recently, several studies have integrated intermediate reasoning process into GR (Figure 1(a)), generating item semantic tokens to construct paths toward final predictions and achieving significant performance gains [10, 3, 4]. However, existing models face several limitations. First, each item is deterministically encoded into a fixed sequence of semantic tokens, producing rigid representations that conflict with the inherent uncertainty of reasoning. Since inference paths are constructed from these fixed item representations, the singularity of the representations directly results in a lack of diversity in the exploration space of reasoning pathways. Items have multi-dimensional attributes such as brand, style, and color, while different users may have diverse reasons for interest in the same item. Consequently, the constrained reasoning space of existing models fails to adequately capture this diversity of user intents, ultimately leading to suboptimal recommendation performance. Second, current methods lack mechanisms to assess or calibrate the reliability of reasoning paths, which may introduce uncontrolled noise during the reasoning process and reduce recommendation accuracy. To overcome these limitations, we aim to develop a reliable GR reasoning framework that supports diverse reasoning paths. To this end, we propose two research problems:
  • (RQ1) How can we design flexible reasoning paths to accommodate diverse user intents?
  • (RQ2) How can we ensure a reliable reasoning process to enhance recommendation quality?
@@
Figure 1: Comparison of reasoning in GR: (a) GR with a limited reasoning process; (b) our REG4Rec enables diverse and reliable reasoning process to better capture the variety of users’ intents.
@@
To address these questions, we draw on core mechanisms of Large Language Models (LLMs). Their success in complex tasks stems from two advances: adopting dynamic, multi-step reasoning beyond linear thought, and incorporating self-reflective critique to improve output reliability [33, 34, 19]. Building on these ideas, we introduce REG4Rec, a reasoning-enhanced generative recommendation model (Figure 1 (b)). To design flexible reasoning paths (RQ1), we introduce the Mixture-of-Experts (MoE)–based Parallel Quantization Codebooks (MPQ) for each item. Given multimodal inputs, MPQ uses parallel autoencoder experts to generate distinct, unordered semantic codebooks, each representing a different facet of the item. At each step, REG4Rec generates parallel predictions from all codebooks and dynamically selects the token with the highest confidence, enabling a vast combinatorial variety of paths. To ensure a reliable reasoning process (RQ2), we propose a comprehensive framework for training and inference. Specifically, we design a Preference Alignment Reward for Selection (PARS) using reinforcement learning (RL) [22] to guide the model in selecting inference paths with high consistency and confidence. Furthermore, we propose Multi-Step Reward Augmentation (MSRA), which extends the reward horizon to future user actions, enabling better capture of long-term preferences and reducing noise from stochastic user behavior. During inference, we employ Consistency-Oriented Self-Reflection for Pruning (CORP) to identify and correct or discard inconsistent reasoning paths, thereby mitigating noise and preventing the propagation of erroneous reasoning. To make these post‑training refinements practical at industrial scale, we design an online Layer‑Adaptive Dynamic Quantization Controller (LADQ) that periodically profiles per‑layer sensitivity and assigns fp32/bf16/fp8 precision to minimize training time under an accuracy budget.
@@
We conduct extensive experiments on four datasets (three public datasets and a large-scale industrial dataset from an online advertising system), comparing REG4Rec with six state-of-the-art baselines, such as GR models with reasoning process like ReaRec and STREAM. REG4Rec consistently outperforms all baselines across every scenario, achieving performance gains of up to 16.59%. To validate its real-world effectiveness, we deploy REG4Rec on Alibaba’s large-scale commercial advertising platform. The subsequent online A/B test yielded a 5.60% increase in Advertising Revenue, a 1.81% increase in Click-Through Rate (CTR), and a 3.29% increase in Gross Merchandise Volume (GMV). These results demonstrate the significant practical value and real-world impact of our model.
@@
II Related Work
II-A Sequential Recommendation
Sequential recommendation is a fundamental task widely adopted in search, recommendation, and advertising systems [10, 1, 5]. It predicts a user’s next item based on their chronological interaction history. To uncover latent behavioral patterns, researchers have developed various sequential modeling methods, including CNNs [14, 15], RNNs [16, 17], and Transformers [2, 6, 7]. For instance, SASRec [2] and BERT4Rec[6] demonstrate that Transformer architectures effectively track changes in users’ interests, while PinnerFormer [7] further extends this approach to model long-term preferences. While these models are effective, they often rely heavily on item identifiers. In large-scale recommendation scenarios, the presence of billions of distinct items can worsen the Matthew effect, resulting in poorly trained identifier parameters and restricted scalability [13]. To address this, several works have replaced massive item-ID vocabularies with compact semantic tokens in GR models. For instance, Google’s TIGER [18] employs a residual quantized variational autoencoder (RQ-VAE) to create semantic tokens instead of item IDs, these tokens are then fed into a GR model to predict subsequent token IDs. Similarly, OneRec [1] uses a residual k-means codebook to encode items and generate recommendation lists through preference alignment, while LIGER [24] and COBRA [10] leverages a multi-granularity GR framework that integrates semantic tokens with sparse item identifiers.
@@ II-A Sequential Recommendation
Existing GR models, which generate recommendations in a single step, often overlook the reasoning needed to capture complex intents. In contrast, we propose a reasoning-enhanced model that adaptively selects reliable reasoning paths, delivering more accurate and insightful recommendations.
II-B Reasoning-Enhanced Generation
@@
Reasoning is fundamental to solving complex problems [25, 26]. Recent breakthroughs in LLMs have moved beyond single-step generation, driven by two key advancements. First, LLMs create adaptive, structured reasoning flows. Instead of following a single linear path, models now explore a dynamic space of possibilities, either internally through branching thought-paths (e.g., Tree of Thoughts [34]) or externally by interacting with tools (e.g., ReAct [33]). Second, they integrate reflection mechanisms for self-critique [19, 20]. This involves a model drafting, evaluating, and iteratively correcting its own solutions, which significantly enhances output reliability. Inspired by these LLM advancements, researchers have begun to incorporate reasoning into recommendation systems [3, 4]. However, current approaches often implement these principles in a limited fashion. For instance, while ReaRec [3] generates intermediate outcomes, its implicit nature lacks the verifiable structure required for robust reasoning. Similarly, although STREAM [4] constructs explicit reasoning paths, its approach is overly rigid: it confines paths to textually similar items and represents each item with a fixed sequence of semantic tokens. This rigidity prevents the model from capturing the diverse intents users may have for the same item, which necessitates more adaptive reasoning.
@@
Figure 2: REG4Rec introduces MPQ for generating flexible and reliable reasoning paths. During training, CRSS diversifies path combinations, while PARS selects consistent, confident paths. MSRA extends the reward horizon to capture long-term preferences and reduce noise caused by stochastic behavior. During inference, CORP prunes inconsistent paths, preventing the propagation of errors in the reasoning process.
III Methodology
III-A Problem Formulation
This section formalizes the GR task. Let $\mathcal{I}$ denote the set of all items. For a given user $u$, the historical interaction sequence is defined as $\mathcal{S}_{u}=(i_{1},i_{2},\ldots,i_{T})$, where $i_{t}\in\mathcal{I}$ is the item interacted with at time step $t$. The primary objective is to predict the user’s next interaction, $i_{T+1}\in\mathcal{I}$. Inspired by prior work [18, 10], we integrate the GR model into the retrieval stage of recommender systems. A common approach is to build an encoder-decoder GR model. The model takes $\mathcal{S}_{u}$ as input and estimates a distribution $p(i|\mathcal{S}_{u})$ over $i\in\mathcal{I}$. During inference, the model retrieves the top-$N$ items by sorting $p(i|\mathcal{S}_{u})$ in descending order.
III-B Overview of REG4Rec
@@
Our method addresses three key objectives:
  • (i) designing flexible reasoning steps to accommodate diverse user intents (Section IV).
  • (ii) enhancing reasoning reliability to improve recommendation performance (Sections V and VI-B).
  • (iii) introducing an efficient offline training strategy for large-scale recommendation (Section VI-A).
IV Flexible Reasoning Paths
@@
IV-1 MPQ: Moe-based Parallel Quantization Codebooks
At the item tokenization stage, in contrast to traditional RQ codebooks [10], we propose the MPQ for the reasoning process in REG4Rec. MPQ uses multiple codebooks in parallel, which are order-invariant. Each codebook captures a distinct semantic dimension of an item, yielding a factorized representation without a predefined hierarchy. Leveraging an MoE-style gating mechanism [23], MPQ enables information sharing across codebooks and adaptively adjusts their granularity and diversity to better capture semantics. As a result, MPQ balances the information used at each reasoning step with the overall size of the reasoning space.
Concretely, for an item $i$, its multi-modal features (e.g., textual, visual, and behavioral) are extracted with a pretrained model [21] and concatenated into a unified feature vector $v_{i}$. Let $M$ denote the number of parallel autoencoder experts $\{E_{r}\}^{M}_{r=1}$. Each expert $E_{r}$ is associated with a codebook $\mathcal{C}^{r}=\{z^{r}_{1},z^{r}_{2},\ldots,z^{r}_{D}\}$, where $z^{r}_{j}$ is a codeword and $D$ is the codebook size. Each expert operates independently, takes $v_{i}$ as input and projects it into a latent spaces $e^{r}_{i}=E_{r}(v_{i})$. The experts then select the most similar codeword in $\mathcal{C}_{r}$ to $e^{r}_{i}$, using its index to represent a semantic token for item $i$. Formally, the index $c^{i}_{r}$ which is selected by the $r$-th expert represent the semantic token, is calculated as:
$c^{i}_{r}=\arg\min_{j}\|e^{r}_{i}-z^{r}_{j}\|_{2},$ (1)
where $\arg\min$ returns the index of the codeword with the smallest L2 distance [41] to $e^{r}_{i}$. After obtaining $z^{r}_{i}$ from each expert, we employ a routing gate to aggregate the semantic contributions from all codebooks. The final encoding vector is $\mathbf{q}_{i}=\sum_{r=1}^{M}h^{r}(v_{i})z^{r}_{i}$, where $h^{r}(v_{i})$ denotes the routing weight assigned to expert $r$, based on the input feature vector $v_{i}$.